# Denoising Diffusion Variational Inference: Diffusion Models as Expressive Variational Posteriors

### v3

@@ 3.4 Combining The Wake-Sleep Regularized ELBO With Diffusion Models
where $d$ is the dimension of $\mathbf{y}$ and we use the notation $\mathbf{y}_{T+1}=\mathbf{x}$. The right-hand term can be approximated using Monte Carlo; it is also common to leave the variance $\Sigma_{\bm{\phi}}$ fixed (as we typically do in our experiments), in which case $H(q)$ is a constant.
@@ 4.1 Semi-Supervised Learning
Following Makhzani et al. [20], we extend our algorithm to the semi-supervised learning setting where some data points have labels denoted by $l$. We assume the user provides a model of the form $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z},l)=p_{\bm{\theta}}(\mathbf{x}|\mathbf{z},l)r(\mathbf{y}|\mathbf{z},l)p_{\bm{\theta}}(\mathbf{z}|l)p(l)$; we set the variational distributions to $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x},\mathbf{y},l),q_{\bm{\phi}}(\mathbf{y}|\mathbf{x}),q_{\bm{\phi}}(l|\mathbf{x})$. In this setting, we consider two cases, depending on whether the label is observed [15]. We extend Equation 8 to incorporate the label $l$ corresponding to a data point as follows:
$\mathcal{L}_{\text{semi}}=$ $~{}\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x},l)}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z},l)]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x},l)||p_{\bm{\theta}}(\mathbf{y},\mathbf{z}|l))$ (14)
$-\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x})}\left[D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x},l)||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x},l))\right]$
When the label $c$ cannot be observed, we treat it as a latent variable and modify the learning objective $\mathcal{U}_{\text{semi}}=\sum_{c}q_{\bm{\phi}}(l|\mathbf{x})\mathcal{L}_{\text{semi}}(\mathbf{x},l,{\bm{{\bm{\theta}}}},{\bm{\phi}})+D_{\mathrm{KL}}(q_{\bm{\phi}}(l|\mathbf{x})||p(l))$. Therefore, we can conclude a marginal likelihood on our dataset as follows: ~$\tilde{\mathcal{L}}_{\text{semi}}=\sum_{(\mathbf{x},l)\in L}\mathcal{L}_{\text{semi}}(\mathbf{x},l,{\bm{{\bm{\theta}}}},{\bm{{\bm{\phi}}}})+\sum_{\mathbf{x}\in U}\mathcal{U}_{\text{semi}}(\mathbf{x},{\bm{{\bm{\theta}}}},{\bm{{\bm{\phi}}}}).$ where $L$ and $U$ are the sets of data with and without labels, respectively.
We also want to guarantee that all model parameters can be learned in all cases, including $q_{\bm{\phi}}(l|\mathbf{x})$, such that this posterior can be applied as a classifier during inference. Thus, we combine the marginal likelihood with a classification loss to form an extended learning objective: ~$\tilde{\mathcal{L}}_{\text{semi}_{\bm{\alpha}}}=\tilde{\mathcal{L}}_{\text{semi}}+{\bm{\alpha}}\cdot\mathbb{E}_{\tilde{p}(\mathbf{x},l)}\left[-\log q_{\bm{\phi}}(l|\mathbf{x})\right]$
@@ 4.2 Clustering
We have further extended our algorithm to encompass the clustering paradigm. We propose two distinct strategies. In the first approach, we simply fir a model in which $p_{\bm{{\bm{\theta}}}}(\mathbf{z})$ as a mixture of desired priors. The means of these priors are characterized by ${\bm{{\bm{\theta}}}}$. From these means, cluster membership, denoted as $\mathbf{c}$ can be deduced. This approach requires no alteration to the existing learning objective.
Alternatively, the second method retains the original prior but introduces an additional cluster latent variable $\mathbf{c}$ where $\sum_{i}c_{i}=1$. Thus, the model can be specified as $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z},\mathbf{c})=p_{\bm{\theta}}(\mathbf{x}|\mathbf{z},\mathbf{c})r(\mathbf{y}|\mathbf{z})p_{\bm{\theta}}(\mathbf{z})p(\mathbf{c})$ with $p(\mathbf{c})=Dir({\bm{{\bm{\epsilon}}}})$. Consequently, the variational distributions become $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y},\mathbf{c},\mathbf{x}),q_{\bm{\phi}}(\mathbf{y},\mathbf{c}|\mathbf{x})$. This reformulates the learning objective as:
$\mathcal{L}_{\text{clus}}(\mathbf{x})=$ $~{}\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z},\mathbf{c}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z},\mathbf{c})]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z},\mathbf{c}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{y},\mathbf{z},\mathbf{c}))-\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x})}\left[D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))\right]$ (15)
Figure 2: Unsupervised visualization on MNIST using three different priors (pinwheel, swiss roll, and square). Each color indicates a class.
Method Pinwheel Swiss Roll Square ELBO MMD Latent NLL ELBO MMD Latent NLL ELBO MMD Latent NLL AEVB $-12.13\pm 0.41$ $0.77\pm 0.04$ $1.68\pm 0.31$ $-14.80\pm 0.23$ $0.78\pm 0.17$ $5.65\pm 1.58$ $-7.85\pm 0.29$ $1.10\pm 0.66$ $2.78\pm 0.61$ AEVB-IAF $-4.19\pm 0.05$ ${0.77\pm 0.00}$ $1.64\pm 0.73$ $-5.10\pm 0.30$ $0.61\pm 0.15$ $4.43\pm 1.09$ $-3.97\pm 0.22$ $0.75\pm 0.12$ $1.68\pm 0.27$ AAEB $N/A$ $0.68\pm 0.02$ $1.54\pm 0.19$ $N/A$ $\bm{0.52\pm 0.03}$ ${3.34\pm 0.16}$ $N/A$ $0.80\pm 0.02$ $2.46\pm 0.46$ H-AEVB $-7.03\pm 3.13$ ${0.74\pm 0.02}$ $2.25\pm 3.02$ $-7.21\pm 4.62$ ${0.70\pm 0.22}$ $4.04\pm 4.62$ $-5.71\pm 3.05$ ${0.76\pm 0.21}$ ${2.22\pm 2.03}$ PIS $-7.83\pm 0.64$ $0.75\pm 0.14$ $6.50\pm 1.11$ $-9.83\pm 0.61$ $0.61\pm 0.03$ $\bm{2.40\pm 1.01}$ $-7.06\pm 0.06$ $0.77\pm 0.04$ $3.67\pm 0.08$ DDVI $\bm{-3.88\pm 0.96}$ $\bm{0.67\pm 0.04}$ $\bm{1.27\pm 0.21}$ $\bm{-5.03\pm 0.58}$ $0.62\pm 0.33$ $3.86\pm 0.17$ $\bm{-3.79\pm 0.14}$ $\bm{0.66\pm 0.07}$ $\bm{1.56\pm 0.09}$

### v4

@@ 3.2 Learning Objective: A Markovian ELBO
where $d$ is the dimension of $\mathbf{y}$ and we use the notation $\mathbf{y}_{T+1}=\mathbf{x}$. It is also common to leave the variance $\Sigma_{\boldsymbol{\phi}}$ fixed, in which case $H(q)$ is a constant.
3.3 Refining the Objective: A Regularized ELBO
Notice that optimizing $\mathcal{L}_{\text{reg}}$ involves sampling from the approximate reverse process $q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$ to match the true reverse process $r(\mathbf{y}|\mathbf{z},\mathbf{x})$: this is the opposite of diffusion training, where we would sample from $r$ to fit $q$. This type of on-policy learning of $q$ has been studied in the context of approximate inference (Zhang and Chen 2021); however, it requires backpropagating through $T$ samples, which may hamper training, and it optimizes a mode-covering divergence that may struggle to fit complex $p(\mathbf{z})$.
Adding Wake-Sleep Regularization to the ELBO
We propose to further improve the ELBO via off-policy diffusion-like training. Our new objective is the ELBO in Equation (5) augmented with a regularizer $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$.
$\log p_{\boldsymbol{\theta}}(\mathbf{x})\geq$ $\underbrace{\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})]}_{\text{wake / recons. term $\mathcal{L}_{\text{rec}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$}}$ (7)
$\underbrace{-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))}_{\text{priorregularization term $\mathcal{L}_{\text{reg}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$}}$
$\underbrace{-\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x})}[\text{KL}(p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})||q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x}))]}_{\text{sleep term $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$}}$
The optimization of the regularizer $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$ is similar to the sleep phase of wake-sleep, and closely resembles diffusion model training (see below). As in wake-sleep, $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$ is optimized over ${\boldsymbol{\phi}}$ only, the $\mathbf{x}$ are sampled from $p$.
From Wake-Sleep to Diffusion Regularization
Computing $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$ still involves intractable distributions $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}),q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$. To optimize $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$, we introduce another lower bound $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$, which we call the denoising diffusion loss (for reasons that will become apparent shortly):
$\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$ $=-\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x})}[\text{KL}(p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})||q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x}))]$ (8)
$=\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})}[\log{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}]+\bar{H}(p_{\boldsymbol{\theta}})$
$\geq\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})}\left[\mathbb{E}_{r}[\log\frac{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}{r(\mathbf{y}|\mathbf{z},\mathbf{x})}]\right]+\bar{H}(p_{\boldsymbol{\theta}})$
$=\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$
In Equation (8), we applied the ELBO with $r(\mathbf{y}|\mathbf{z},\mathbf{x})$ playing the role of the variational posterior over the latent $\mathbf{y}$ in $q_{\boldsymbol{\phi}}$; $\bar{H}(p_{\boldsymbol{\theta}})$ is the expected conditional entropy of $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$, a constant that does not depend on ${\boldsymbol{\phi}}$.
We can further simplify $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$ by leveraging the Markov structure of the forward and reverse processes $r,q$. Recall that each $\mathbf{y}=(\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{T})$ can be a vector of $T$ latents, which we also denote as $\mathbf{y}_{1:T}$, and that $r(\mathbf{y}_{1:T}|\mathbf{z},\mathbf{x})=\prod_{t=1}^{T}r(\mathbf{y}_{t}|\mathbf{y}_{t-1},\mathbf{x})$, where $\mathbf{y}_{0}=\mathbf{z}$ and also $q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})=q_{\boldsymbol{\phi}}(\mathbf{y}_{0:T}|\mathbf{x})=q_{\boldsymbol{\phi}}(\mathbf{y}_{T}|\mathbf{x})\prod_{t=1}^{T}q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x}).$
We may use the Markov structure in $q,r$ to rewrite $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$ as a sum of $T$ terms, one per Markov step. The derivation is identical to that used to obtain the ELBO of a diffusion model (Sohl-Dickstein et al. 2015), and yields an expression of the same form:
$\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})=$ $~{}\mathbb{E}_{r}\left[\log q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})-\sum_{t=2}^{T}\text{KL}(r_{t}||q_{t})\right]$ (9)
$-\text{KL}(r(\mathbf{y}_{T}|\mathbf{z})||q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x})).$
where $r_{t},q_{t}$ denote the distributions $r(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{y}_{0},\mathbf{x})$ and $q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})$ (see Appendix K for the derivation).
Regularized DDVI Objective
We define the full DDVI objective $\mathcal{L}_{\text{ddvi}}$ to be the sum of the aforementioned terms:
$\mathcal{L}_{\text{ddvi}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})=\mathcal{L}_{\text{rec}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})+\mathcal{L}_{\text{reg}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})+\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$
Terms $\mathcal{L}_{\text{reg}}$ and $\mathcal{L}_{\text{diff}}$ may be weighted by hyper-parameters $\beta_{\text{reg}},\beta_{\text{diff}}>0$, as in the $\beta$-VAE framework. In our experiments, $\beta_{\text{reg}}=\beta_{\text{diff}}=1$ unless otherwise specified. Note that since $\mathcal{L}_{\text{diff}}\leq\mathcal{L}_{\text{sleep}}\leq 0$, $\mathcal{L}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ is a valid lower bound on $\log p_{\boldsymbol{\theta}}(\mathbf{x})$ that is tight when $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})=p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$.
3.4 Optimization: Extending Wake-Sleep
We may optimize $\mathcal{L}_{\text{ddvi}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ using gradient descent by alternating between ELBO optimization and taking sleep steps (see Appendix A for full details):
  1. 1.
    Sample $\mathbf{x}$ from data, take gradient step on ${\boldsymbol{\theta}},{\boldsymbol{\phi}}$ optimizing $\mathcal{L}_{\text{rec}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})+\mathcal{L}_{\text{reg}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ (the “wake” step);
  2. 2.
    Sample $\mathbf{z},\mathbf{x}$ from $p_{\boldsymbol{\theta}}$ and take a gradient step on ${\boldsymbol{\phi}}$ optimizing $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$ (the “sleep” step).
Again, terms may be weighted by $\beta_{\text{reg}},\beta_{\text{diff}}>0$. Note that by maximizing $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$, we fit $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ to $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$ via the forward KL divergence; similarly, by optimizing $\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{reg}}$ (the ELBO), we fit $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ to $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$ via the reverse KL divergence. Thus, optimizing $\mathcal{L}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ encourages $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ to approximate $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$, and when the two are equal, the bound $\mathcal{L}_{\text{ddvi}}$ on $\log p_{\boldsymbol{\theta}}(\mathbf{x})$ is tight.
Simplified Wake-Sleep
We also consider a light-weight algorithm, in which $r(\mathbf{y}|\mathbf{z})$ and $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y})$ do not depend on $\mathbf{x}$. This scenario admits the following optimization procedure:
  1. 1.
    Sample $\mathbf{x}$ from data and compute gradient on ${\boldsymbol{\theta}},{\boldsymbol{\phi}}$ optimizing $\mathcal{L}_{\text{rec}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})+\mathcal{L}_{\text{reg}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$.
  2. 2.
    Sample $\mathbf{z}$ from $p(\mathbf{z})$ and compute gradient on ${\boldsymbol{\phi}}$ optimizing $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$; take step on weighted sum of both gradients.
In this case, $\mathcal{L}_{\text{diff}}$ requires only sampling from $p(\mathbf{z})$, and the entire loss $\mathcal{L}_{\text{ddvi}}$ can be optimized end-to-end using gradient descent. This algorithm is simpler (there is no separate sleep phase); however, $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ may not perfectly approximate $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$ when $r(\mathbf{y}|\mathbf{z})$ and $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y})$ do not depend on $\mathbf{x}$, hence $\mathcal{L}$ may no longer be a tight bound.
Practical Considerations
A common type of noising process compatible with this bound when $\mathbf{z}$ is continuous is Gaussian diffusion, where we define $r(\mathbf{y}_{t}|\mathbf{y}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\sqrt{1-{\boldsymbol{\alpha}}_{t}}\mathbf{y}_{t-1},{\boldsymbol{\alpha}}_{t}\mathbf{I})$ for a suitable schedule $({\boldsymbol{\alpha}}_{t})_{t=1}^{T}$. We then adopt the parameterization $q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})=\mathcal{N}(\mathbf{y}_{t-1};\mu_{\boldsymbol{\phi}}(\mathbf{y}_{t},\mathbf{x},t),\Sigma_{\boldsymbol{\phi}}(\mathbf{y}_{t},\mathbf{x},t))$. It is then common to parameterize $q_{\boldsymbol{\phi}}$ with a noise prediction network ${\boldsymbol{\epsilon}}_{\boldsymbol{\phi}}$ (Ho, Jain, and Abbeel 2020); the sum of KL divergences can be approximated by $\mathbb{E}_{t,{\boldsymbol{\epsilon}}_{t}\sim r(\mathbf{y}_{0},t)}||{\boldsymbol{\epsilon}}_{t}-{\boldsymbol{\epsilon}}_{\boldsymbol{\phi}}(\sqrt{\bar{{\boldsymbol{\alpha}}_{t}}}\mathbf{y}_{0}+\sqrt{1-\bar{{\boldsymbol{\alpha}}_{t}}}{\boldsymbol{\epsilon}}_{t},\mathbf{x},t)||^{2}$. Other extensions include discrete denoising processes (Austin et al. 2021; Sahoo et al. 2024; Schiff et al. 2024). In the wake-sleep setting, we know both endpoints $\mathbf{y}_{T}\sim q(\cdot|\mathbf{x})$ and $\mathbf{y}_{0}=\mathbf{z}$ of the diffusion process, opening the possibility for applying optimal transport techniques (Cuturi 2013; De Bortoli et al. 2021).
@@ 4.1 Semi-Supervised Learning
Following Makhzani et al. (2015), we extend our algorithm to the semi-supervised learning setting where some data points have labels denoted by $l$. We assume the user provides a model of the form $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z},l)=p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z},l)r(\mathbf{y}|\mathbf{z},l)p_{\boldsymbol{\theta}}(\mathbf{z}|l)p(l)$; we set the variational distributions to $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x},\mathbf{y},l),q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x}),q_{\boldsymbol{\phi}}(l|\mathbf{x})$. In this setting, we consider two cases, depending on whether the label is observed (Kingma et al. 2014). We extend Equation (7) to incorporate the label $l$ corresponding to a data point as follows:
$\mathcal{L}_{\text{semi}}=$ $~{}\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x},l)}[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z},l)]$ (10)
$-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x},l)||p_{\boldsymbol{\theta}}(\mathbf{y},\mathbf{z}|l))$
$-\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x})}\left[\text{KL}(p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x},l)||q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x},l))\right]$
When the label $c$ cannot be observed, we treat it as a latent variable and modify the learning objective $\mathcal{U}_{\text{semi}}=\sum_{c}q_{\boldsymbol{\phi}}(l|\mathbf{x})\mathcal{L}_{\text{semi}}(\mathbf{x},l,{\boldsymbol{{\boldsymbol{\theta}}}},{\boldsymbol{\phi}})+\text{KL}(q_{\boldsymbol{\phi}}(l|\mathbf{x})||p(l))$. Therefore, we can conclude a marginal likelihood on our dataset as follows: ~$\tilde{\mathcal{L}}_{\text{semi}}=\sum_{(\mathbf{x},l)\in L}\mathcal{L}_{\text{semi}}(\mathbf{x},l,{\boldsymbol{{\boldsymbol{\theta}}}},{\boldsymbol{{\boldsymbol{\phi}}}})+\sum_{\mathbf{x}\in U}\mathcal{U}_{\text{semi}}(\mathbf{x},{\boldsymbol{{\boldsymbol{\theta}}}},{\boldsymbol{{\boldsymbol{\phi}}}}).$ where $L$ and $U$ are the sets of data with and without labels, respectively.
We also want to guarantee that all model parameters can be learned in all cases, including $q_{\boldsymbol{\phi}}(l|\mathbf{x})$, such that this posterior can be applied as a classifier during inference. Thus, we combine the marginal likelihood with a classification loss to form an extended learning objective: ~$\tilde{\mathcal{L}}_{\text{semi}_{\boldsymbol{\alpha}}}=\tilde{\mathcal{L}}_{\text{semi}}+{\boldsymbol{\alpha}}\cdot\mathbb{E}_{\tilde{p}(\mathbf{x},l)}\left[-\log q_{\boldsymbol{\phi}}(l|\mathbf{x})\right]$
@@ 4.2 Clustering
We have further extended our algorithm to encompass the clustering paradigm. We propose two distinct strategies. In the first approach, we simply formulate a model in which $p_{\boldsymbol{{\boldsymbol{\theta}}}}(\mathbf{z})$ is a mixture of desired priors. The means of these priors are characterized by ${\boldsymbol{{\boldsymbol{\theta}}}}$. From these means, cluster membership, denoted as $\mathbf{c}$ can be deduced. This approach requires no alteration to the existing learning objective.
Alternatively, the second method retains the original prior, but introduces an additional latent cluster variable $\mathbf{c}$ where $\sum_{i}c_{i}=1$. Thus, the model can be specified as $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z},\mathbf{c})=p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z},\mathbf{c})r(\mathbf{y}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})p(\mathbf{c})$ with $p(\mathbf{c})=Dir({\boldsymbol{{\boldsymbol{\epsilon}}}})$. Consequently, the variational distributions become $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y},\mathbf{c},\mathbf{x}),q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{c}|\mathbf{x})$. This yields the objective:
$\mathcal{L}_{\text{clus}}(\mathbf{x})=$ $~{}\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z},\mathbf{c}|\mathbf{x})}[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z},\mathbf{c})]$ (11)
$-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z},\mathbf{c}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{y},\mathbf{z},\mathbf{c}))$
$-\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x})}\left[\text{KL}(p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})||q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x}))\right]$
Expectations over small numbers of classes $\mathbf{c}$ are done analytically; larger $\mathbf{c}$ require backpropagating through discrete sampling (Jang, Gu, and Poole 2016; Sahoo et al. 2023).
Figure 2: Unsupervised visualization on MNIST using three priors (pinwheel, swiss roll, and square). Each color indicates a class.
Method Pinwheel Swiss Roll Square ELBO MMD Latent NLL ELBO MMD Latent NLL ELBO MMD Latent NLL AEVB $-12.13\pm 0.41$ $0.77\pm 0.04$ $1.68\pm 0.31$ $-14.80\pm 0.23$ $0.78\pm 0.17$ $5.65\pm 1.58$ $-7.85\pm 0.29$ $1.10\pm 0.66$ $2.78\pm 0.61$ AEVB-IAF $-4.19\pm 0.05$ ${0.77\pm 0.00}$ $1.64\pm 0.73$ $-5.10\pm 0.30$ $0.61\pm 0.15$ $4.43\pm 1.09$ $-3.97\pm 0.22$ $0.75\pm 0.12$ $1.68\pm 0.27$ AAEB $N/A$ $0.68\pm 0.02$ $1.54\pm 0.19$ $N/A$ $\mathbf{0.52\pm 0.03}$ ${3.34\pm 0.16}$ $N/A$ $0.80\pm 0.02$ $2.46\pm 0.46$ H-AEVB $-7.03\pm 3.13$ ${0.74\pm 0.02}$ $2.25\pm 3.02$ $-7.21\pm 4.62$ ${0.70\pm 0.22}$ $4.04\pm 4.62$ $-5.71\pm 3.05$ ${0.76\pm 0.21}$ ${2.22\pm 2.03}$ PIS $-7.83\pm 0.64$ $0.75\pm 0.14$ $6.50\pm 1.11$ $-9.83\pm 0.61$ $0.61\pm 0.03$ $\mathbf{2.40\pm 1.01}$ $-7.06\pm 0.06$ $0.77\pm 0.04$ $3.67\pm 0.08$ DDVI $\mathbf{-3.88\pm 0.96}$ $\mathbf{0.67\pm 0.04}$ $\mathbf{1.27\pm 0.21}$ $\mathbf{-5.03\pm 0.58}$ $0.62\pm 0.33$ $3.86\pm 0.17$ $\mathbf{-3.79\pm 0.14}$ $\mathbf{0.66\pm 0.07}$ $\mathbf{1.56\pm 0.09}$